Coline has the following capital gain and loss transactions for 2018. a. Short-term capital gain b. Short-term capital loss Long-term capital gain (28%) Long-term capital gain (15%) Long-term capital loss (28%) $10,700 4,494) 12,840 4,280 (22,470) d. After the netting process, Coline has an overall net short-term capital gain of $ 4,151 X
Answer #1

Net STCG = 10,700 - 4,494 = 6206 STCG

Net LTCL = 12,840 + 4,280 - 22470 = 5350 LTCL

After netting

= 6206 STCG - 5350 LTCL

= 856 STCG

= Net short tem capital gain of 856